# Dissipating quartets of excitations in a superconducting circuit

### v2

@@ Acknowledgments
Figure S13: Simplified six-level model employed to fit the datasets presented in Fig. 4. The pump (plain purple arrow) and the linear drive applied to the buffer (plain orange arrow) are each resonant with a single transition only (detuned transitions represented by dashed arrows). As a result, the system state is confined to a six-level manifold (levels pictured in black) and never visits other states (levels pictured in gray).
We first notice that for both datasets presented Fig. 3b-c, the system state is constrained to a 6-dimension manifold spanned by $\{|0\rangle_{a}|0\rangle_{b},|1\rangle_{a}|0\rangle_{b},|2\rangle_{a}|0\rangle_{b},|3\rangle_{a}|0\rangle_{b},|4\rangle_{a}|0\rangle_{b},|0\rangle_{a}|1\rangle_{b}\}$ (see Fig. S13). Indeed, the pump (purple arrow) is only resonant with the $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition and is weak enough that no other transition can be driven ($g_{4}\sqrt{\frac{(n+4)!}{n!}(m+1)}\ll|(\chi_{n+4,m}-\chi_{n,m+1})|$ for $n>0$ or $m>0$). Similarly, the linear drive on the buffer (only on when recording the reflection spectra in Fig. 3c) is resonant with the $|0\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition only ($\xi_{b}\sqrt{m+1}\ll|(\chi_{n,m+1}-\chi_{n,m})|$ for $n>0$ or $m>0$).
@@
For the data presented in Fig. 3c, reduction of the system Hilbert space is less obvious. Here, we are interested in the steady-state value of $\langle{\bf b}\rangle$. We directly apply the Lindblad adjoint formalism to encode the evolution of $\langle{\bf b}\rangle$ along with the expectation value of ten other orthogonal operators (which we compute numerically and do not detail here) in a $11\times 1$ 11 matrix $M$. The steady-state is directly found from the single vector in the kernel of $M$.
With these model simplifications, we perform a global fit of the two datasets in Fig. 3b-c with a simple least-square method. As a sanity check of our simplifying assumptions, we finally perform a full Lindblad master equation featuring the effective Hamiltonian ${\bf H}_{RWA}$ and the dissipators $\kappa_{b}\mathcal{D}[{\bf b}]$, $2\kappa_{\phi}\mathcal{D}[{\bf a^{\dagger}}{\bf a}]$ and $\gamma_{k}\mathcal{D}[|k-1\rangle\langle k|]$ for $1\leq i\leq 4$, with parameter values extracted from the above fit. The results of these simulations are plotted in Fig. 3b-c and Fig. S10 and quantitatively reproduce our data.
S5 Tuning the system away from TLS resonances
In our experiment, some of the memory transitions—either single-photon or three-photon transitions—were transiently observed to shift or split. Typically, a given transition that would yield a clean Lorentzian response in two-tone spectroscopy would suddenly yield a two-peak response. This behavior was either be permanent (until the system was cycled around 20 K before being cooled back down at base temperature) or appearing on and off on a minute timescale (see Fig. S14). The buffer lowest energy transition was also witnessed to split at one occasion.
@@
Various mechanisms can contribute to a transition frequency shift proportional to the power of an off-resonant drive. In particular, a flux drive applied to the ATS loops with both a symmetric and an anti-symmetric component yields such a shift (see Sec. S2). The Stark effect [88] is another possible mechanism owing to the modes non-linearity and given that their states are linearly displaced by the drive. Such a displacement is expected both for a symmetric flux drive (through the leading term in the ATS sine Hamiltonian) and for an anti-symmetric flux drive (through induction in the circuit loops [85]). This latter mechanism is also expected to degrade the system coherence through measurement-induced dephasing [88]. Measurement-induced transitions are also a concern for some specific drive frequencies [52, 53, 54, 55]. While we did not quantitatively investigate these mechanisms in our experiment, the fact that a frequency shift of the circuit transitions is expected at any flux drive frequency lets us select the one which minimizes decoherence.
@@
In detail, we use the characteristic decay time of Rabi oscillations driven on the memory $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition as a proxy for the system coherence. As illustrated in Fig. S15, we sweep the frequency of an off-resonant drive over a few GHz range. For each frequency, we adjust the drive power to get a 17 MHz shift of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition (which is approximately the value used in our experiment) and record Rabi oscillations in these conditions. We then select the off-resonant drive frequency yielding the longest-lived Rabi oscillations (at 9.25 GHz).
When acquiring the datasets presented in Fig. 3b-c, we apply this off-resonant drive at a power of 5.3 dBm, referenced as in Fig. S15 (at the output of the microwave source labeled Stark shifting pump in Fig. S3). The drive is turned on during the system preparation, application of the four-photon dissipation channel and the measurement. In these conditions, the $T_{1}$ of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition is measured to be of 7.5 $\mu$s (see Fig. S9 for a null pump amplitude) and in a later measurement, we observed the Ramsey free-induction decay time to be of 3.5 $\mu$s, corresponding to a pure dephasing rate of $2\pi\times 35$ kHz. This is in good agreement with the value fitted from the datasets presented in Fig. 3b-c. These lifetimes are slightly degraded when compared to those measured in absence of the off-resonant drive ($T_{1}\lesssim 11~{}\mu$s, $T_{2}\lesssim 6~{}\mu$s), even though these were seen to fluctuate over the course of days.
@@
Figure S14: Signature of hybridization to a Two-Level System. Response of the system when probed in two-tone spectroscopy around the resonance frequency of the $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ transition (driven through the flux lines). The two traces are taken 8 minutes apart. On the later one (orange), the line is split in two peaks of approximately equal height.
@@
Figure S15: Shifting the system transitions out of resonance from TLS’s. A controlled shift of all transitions is realized by the application of a strong off-resonant drive (labeled Stark shifting pump in Fig. S3). In order to select the drive frequency and power minimizing decoherence, we sweep the drive frequency and adjust its power to reach a Stark shift of 17 MHz for the memory $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition (corresponding power referenced at the microwave source output port reported in the bottom panel). We then record Rabi oscillations while this off-resonant drive is on (top panel), and select the drive frequency for which the oscillations are longest-lived (at 9.25 GHz).

### v3

@@ D.2 Dynamics under the four-to one photon exchange interaction
Figure 17: Simplified six-level model employed to fit the datasets presented in Fig. 4. The pump (plain purple arrow) and the linear drive applied to the buffer (plain orange arrow) are each resonant with a single transition only (detuned transitions represented by dashed arrows). As a result, the system state is confined to a six-level manifold (levels pictured in black) and never visits other states (levels pictured in gray).
We first notice that for both datasets presented Fig. 3b-c, the system state is constrained to a 6-dimension manifold spanned by $\{|0\rangle_{a}|0\rangle_{b},|1\rangle_{a}|0\rangle_{b},|2\rangle_{a}|0\rangle_{b},|3\rangle_{a}|0\rangle_{b},|4\rangle_{a}|0\rangle_{b},|0\rangle_{a}|1\rangle_{b}\}$ (see Fig. 17). Indeed, the pump (purple arrow) is only resonant with the $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition and is weak enough that no other transition can be driven ($g_{4}\sqrt{\frac{(n+4)!}{n!}(m+1)}\ll|(\chi_{n+4,m}-\chi_{n,m+1})|$ for $n>0$ or $m>0$). Similarly, the linear drive on the buffer (only on when recording the reflection spectra in Fig. 3c) is resonant with the $|0\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition only ($\xi_{b}\sqrt{m+1}\ll|(\chi_{n,m+1}-\chi_{n,m})|$ for $n>0$ or $m>0$).
@@
For the data presented in Fig. 3c, reduction of the system Hilbert space is less obvious. Here, we are interested in the steady-state value of $\langle{\bf b}\rangle$. We directly apply the Lindblad adjoint formalism to encode the evolution of $\langle{\bf b}\rangle$ along with the expectation value of ten other orthogonal operators (which we compute numerically and do not detail here) in a $11\times 11$ matrix $M$. The steady-state is directly found from the single vector in the kernel of $M$.
With these model simplifications, we perform a global fit of the two datasets in Fig. 3b-c with a simple least-square method. As a sanity check of our simplifying assumptions, we finally perform a full Lindblad master equation featuring the effective Hamiltonian ${\bf H}_{RWA}$ and the dissipators $\kappa_{b}\mathcal{D}[{\bf b}]$, $2\kappa_{\phi}\mathcal{D}[{\bf a^{\dagger}}{\bf a}]$ and $\gamma_{k}\mathcal{D}[|k-1\rangle\langle k|]$ for $1\leq i\leq 4$, with parameter values extracted from the above fit. The results of these simulations are plotted in Fig. 3b-c and Fig. 14 and quantitatively reproduce our data.
Appendix E Tuning the system away from TLS resonances
In our experiment, some of the memory transitions—either single-photon or three-photon transitions—were transiently observed to shift or split. Typically, a given transition that would yield a clean Lorentzian response in two-tone spectroscopy would suddenly yield a two-peak response. This behavior was either be permanent (until the system was cycled around 20 K before being cooled back down at base temperature) or appearing on and off on a minute timescale (see Fig. 18). The buffer lowest energy transition was also witnessed to split at one occasion.
@@ Appendix E Tuning the system away from TLS resonances
Various mechanisms can contribute to a transition frequency shift proportional to the power of an off-resonant drive. In particular, a flux drive applied to the ATS loops with both a symmetric and an anti-symmetric component yields such a shift (see Sec. B). The Stark effect [14] is another possible mechanism owing to the modes non-linearity and given that their states are linearly displaced by the drive. Such a displacement is expected both for a symmetric flux drive (through the leading term in the ATS sine Hamiltonian) and for an anti-symmetric flux drive (through induction in the circuit loops [86]). This latter mechanism is also expected to degrade the system coherence through measurement-induced dephasing [14]. Measurement-induced transitions are also a concern for some specific drive frequencies [48, 79, 54, 7]. While we did not quantitatively investigate these mechanisms in our experiment, the fact that a frequency shift of the circuit transitions is expected at any flux drive frequency lets us select the one which minimizes decoherence.
@@
In detail, we use the characteristic decay time of Rabi oscillations driven on the memory $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition as a proxy for the system coherence. As illustrated in Fig. 19, we sweep the frequency of an off-resonant drive over a few GHz range. For each frequency, we adjust the drive power to get a 17 MHz shift of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition (which is approximately the value used in our experiment) and record Rabi oscillations in these conditions. We then select the off-resonant drive frequency yielding the longest-lived Rabi oscillations (at 9.25 GHz).
[…]
@@
Figure 18: Signature of hybridization to a Two-Level System. Response of the system when probed in two-tone spectroscopy around the resonance frequency of the $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ transition (driven through the flux lines). The two traces are taken 8 minutes apart. On the later one (orange), the line is split in two peaks of approximately equal height.
@@
Figure 19: Shifting the system transitions out of resonance from TLS’s. A controlled shift of all transitions is realized by the application of a strong off-resonant drive (labeled Stark shifting pump in Fig. 7). In order to select the drive frequency and power minimizing decoherence, we sweep the drive frequency and adjust its power to reach a Stark shift of 17 MHz for the memory $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition (corresponding power referenced at the microwave source output port reported in the bottom panel). We then record Rabi oscillations while this off-resonant drive is on (top panel), and select the drive frequency for which the oscillations are longest-lived (at 9.25 GHz).